To buy a computer system, a customer can choose one of 4 monitors, one of 2 keyboards, one of 4 computers and one of 3 printers. Determine the number of possible systems that a customer can choose from.


1
Expert's answer
2021-02-24T06:07:10-0500

Solution:

Number of monitors available = 4

Number of keyboards available = 2

Number of computers available = 4

Number of printers available = 3

A customer needs to select only one item out of their respective numbers.

The number of possible systems that a customer can choose from are "^4 C_1 \times ^2 C_1\times ^4 C_1\times^3 C_1"

"=4\times2\times4\times3" ["\because\ ^n C_1=n" ]

"=96"

Thus, final answer is 96.